OpenStudy (anonymous):

can any one say integration of sin(x)/x in the limit of -infinity to +infinity

OpenStudy (anonymous):

\[\int\limits _{- \infty}^{\infty} \frac{\sin(x)}{x} dx\] In order for you to find the integration of this one, you must break your integral into 2 parts:\[\int\limits_{- \infty}^{0} \frac{\sin(x)}{x} dx + \int\limits_{0}^{\infty}\frac{\sin(x)}{x} dx\] then you can cmpute the integral :)